# Sample complexity of optimal transport barycenters with discrete support

### v2

@@ 5.3.2 Dependency in $N$ in the estimation error
The lower bound (5.2) of the estimation error is of order $O(N^{-\frac{2}{d}}\sqrt{N/n})$ whereas the upper bound (4.2) is of order $O(\sqrt{N/n})$. Thus, this raises the question of the optimal exponent in $N$, which we investigate in this section by exhibiting a situation where the lower bound is tight.
Let us consider a particular case: $N=n$, $L=1$ and $\pi=(1/n,\ldots,1/n)$, the vector of uniform weights. In this setting, the upper bound (4.2) of Corollary 4.4 does not vanish as $n$ increases. However, we can show that the estimation error is upper bounded by a quantity of order $O(n^{-2/d})$. Indeed, in this setting, the empirical barycenter, i.e. the minimizer of $\mu\mapsto W_{2}^{2}(\mu_{n},\mu)$ over the space $\mathcal{M}_{1}^{N}(\mathbb{R}^{d})$, is clearly $\mu_{n}$ and thus the estimation error writes
$\mathbb{E}\left[W_{2}^{2}(\mu,\mu_{n})-\min_{\nu\in\mathcal{M}_{1}^{n}(\mathbb{B}_{R})}W_{2}^{2}(\mu,\nu)\right].$
This quantity is upper bounded by $\mathbb{E}\left[W_{2}^{2}(\mu,\mu_{n})\right]$, which is of order $O(n^{-2/d})$, see fournier2015rate. This shows that the upper bound derived in Corollary 4.4, which scales in this case as $O\big(1\big)$, is not optimal in that case. As suggested by the above bounding argument, finding a better exponent in $N$ is related to estimating convergence rates for $W_{2}(\mu,\mu_{n})$, a notoriously difficult problem in the statistical optimal transport literature (chewi2024statistical, Section 2). This also indicates that the curse of dimensionality is in fact hidden in the sparsity level $N$, an observation already made in the remark following (del2024central, Theorem 2.10).
@@ 5.3.3 Beyond the $W_{2}^{2}$ case
The optimality of the rate of convergence $1/\sqrt{n}$ remains an open question for the other divergences $\mathbb{D}$ which we considered. We expect the same conclusion to hold for $\mathbb{D}=SW_{p}^{p}$ and $\text{max-}SW_{p}^{p}$ since the $1/\sqrt{n}$ rate is also known to be optimal for these divergences xu2022central. Regarding the divergence $\mathbb{D}=W_{\epsilon,p}^{p}$, the empirical Sinkhorn bias benefits from a more favorable $1/n$ convergence rate del2023improved. This observation suggests that a sharper convergence rate might be attainable.
@@ 5.4 K-means and constrained K-means
Similarly to the previous section, choosing $\mathbb{D}=W_{2}^{2}$ and $L=1$ in Corollary 4.4 provides error bounds for the celebrated K-means and constrained K-means problems. Let $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R})$ and $X_{1},\ldots,X_{n}$ be $n$ i.i.d samples of law $\mu$. It is well known that the minimization of the K-means functional
$\min_{Y\in(\mathbb{B}_{R})^{N}}\sum_{i=1}^{n}\min_{j=1,\ldots,N}\|X_{i}-y_{j}\|^{2}$ (5.3)
can be recasted as an optimal quantization problem for the empirical measure $\mu_{n}$ distributed on the sample points $X_{1},\ldots,X_{n}$ (see Appendix D.4). In other words, we have
@@
Corollary 4.4 then yields a convergence rate of order $O(\sqrt{N/n})$ for the K-means problem, which improves upon the convergence rate $O\left(\sqrt{\frac{N\log(N)}{n}}\right)$ given in (gyorfi2002principles, Section 4.4). Additionally, the Wasserstein functional formulation of the K-means problem provides a complementary approach to the K-means problem, closer to the approximation of measure point of view. It also allows to consider arbitrary cluster size constraints, a variant initially formulated in ng2000note. Formally, the constrained K-means problem consists in fixing a vector of weights $\bar{\pi}\in\Delta_{N}$ and then solving
$\min_{Y\in(\mathbb{R}^{d})^{N}}\ W_{2}^{2}\left(\mu_{n},\sum_{i=1}^{N}\bar{\pi}_{i}\delta_{y_{i}}\right).$ (5.5)
For a known target measure $\mu$ and a uniform vector of weights $\bar{\pi}_{i}=1/N$ for all $i\in[\![1,N]\!]$, we recover the optimal uniform quantization problem which is solved with a variation of the Lloyd algorithm nonassymptotic. In the empirical case $\mu_{n}$, Corollary 4.4 applies to this problem as well.
@@ 6 Numerical experiments
In this section, we illustrate  [^1] The code to reproduce the experiments is available in Python at https://github.com/LeoPtls the rate of convergence obtained in Corollary 4.4 for $\mathbb{D}=W_{2}^{2}$ and $W_{\epsilon,2}^{2}$.
Tightness in $n$.
We consider $L=3$ target Gaussian distributions in dimension $d=2$ with random means and random covariance matrices. We compute the $W_{2}^{2}$ and $W_{\epsilon,2}^{2}$ empirical free support barycenter for $N=10$, $N=50$ and for $n$ ranging between $1000$ and $10000$ using the POT library flamary2021pot. The regularization parameter for $W_{\epsilon,2}^{2}$ is set to $\epsilon=1$. Figure 1 displays the evolution of the estimation error in (4.2) as $n$ increases. The observed convergence rate matches the one obtained in Corollary 4.4 for $\mathbb{D}=W_{2}^{2}$ and $\mathbb{D}=W_{\epsilon,2}^{2}$. The presence of hollows in the curve for $\mathbb{D}=W_{\epsilon,p}^{p}$ is due to numerical instabilities.
Figure 1: Estimation error in log-scale as $n$ increases with its associated $95\%$ confidence interval, over $10$ experiments for $\mathbb{D}=W_{2}^{2}$ (top) and $40$ experiments for $\mathbb{D}=W_{\epsilon,2}^{2}$ (bottom). The theoretical bound $n\mapsto 1/\sqrt{n}$ is plotted in orange.
Tightness in $N$.
To illustrate the impact that the sparsity level $N$ has on the estimation error we also consider $L=3$ target Gaussian distributions in dimension $d=6$ with random means and random covariance matrices. We compute the $W_{2}^{2}$ and $W_{\epsilon,2}^{2}$ empirical free support barycenter for fixed sample size $n=500$ and values of $N$ ranging between $10$ and $500$, also using the POT library flamary2021pot. The regularization parameter for $W_{\epsilon,2}^{2}$ is set to $\epsilon=1$. Figure 2 displays the evolution of the estimation error in (4.2). The observed convergence rate lies between the upper bound of Corollary 4.4 and lower bound of Lemma 5.1 for $\mathbb{D}=W_{2}^{2}$, and is close to the upper bound of Corollary 4.4 for $\mathbb{D}=W_{\epsilon,2}^{2}$. This experiment suggests that the dependency in $N$ suggested by the worst case scenario in Lemma 5.1 matches practical observation on mild probability distributions, Gaussian mixtures here.
Figure 2: Estimation error in log-scale as NN increases with its associated $95\%$ confidence interval, over $10$ experiments, and for $n=500$. The theoretical bounds (Corollary 4.4 and Lemma 5.1) are plotted in red.
@@ 7 Conclusion
In this paper, we prove that the rate of convergence in expectation of empirical optimal transport barycenters with sparse support is of order $O\left(\sqrt{\frac{N}{n}}\right)$. This result is described for several types of optimal transport divergences, namely the Wasserstein and sliced distances, as well as the Sinkhorn divergence. We show that they all behaved similarly as the number of observation $n$ per target measure grows to infinity. A natural question is whether these rates are min-max optimal. We have showed that this is true for Wasserstein distances $\mathbb{D}=W_{p}^{p}$. However it remains an open question for sliced distances $\mathbb{D}=SW_{p}^{p},\>\text{max-}SW_{p}^{p}$ and Sinkhorn divergences $\mathbb{D}=W_{\epsilon,p}^{p}$. Finally, although our results are limited to the compactly supported case, we believe this is not a significant restriction, as most applications fall within this setting : texture mixing rabin2011wasserstein, constrained clustering ho2017multilevel, shape interpolation solomon2015convolutional.
@@ Appendix A Semi-dual formulation of optimal transport
The Kantorovich dual formulation of the Wasserstein distances $W_{p}$ can be recast as a semi-dual problem using the $c$-transform. Similarly, the Sinkhorn divergences $W_{\epsilon,p}^{p}$ admit a semi-dual formulation based on the smoothed $c$-transform. In the semi-discrete setting (i.e. transport between a continuous and a discrete measure), these formulations turn an infinite-dimensional program (corresponding to the primal formulations (2.1) and (2.3)) into a finite-dimensional one, which is a key point in our proofs. In this section, we present these semi-dual formulations in the semi-discrete setting, with a particular focus on the case where the discrete measure has zero weights or repeated atoms. Throughout this section, we consider a measure $\mu\in\mathcal{M}_{1}(\mathbb{R}^{d})$.
@@ A.1 Semi-dual formulation outside the generalized diagonal
Let us then define $\nu:=\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}$ a discrete measure supported on $Y:=(y_{1},\ldots,y_{N})\in(\mathbb{R}^{d})^{N}\setminus D_{N}$ and strictly positive weights $\pi\in\mathrm{int}(\Delta_{N})$. For a function bounded continuous function $\phi\in\mathcal{C}_{b}(\mathbb{R}^{d})$ , let us also denote $\phi^{c}\in\mathcal{C}_{b}(\mathbb{R}^{d})$ its c-transform, defined in our case for all $x\in\mathbb{R}^{d}$ as $\phi^{c}(x)=\min_{i=1,...,N}(\|x-y_{i}\|^{p}-\phi(y_{i}))$ (book_santambrogio, Section 1.3). Then by Kantorovich’s duality ((villani_opt_old_new, Theorem 5.10)) and $c$-convexity of the optimal potentials we have the following formulation for the $p$-Wasserstein distance between $\mu$ and $\nu$:
$\begin{split}W_{p}^{p}(\mu,\nu):&=\inf_{\gamma\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{p}d\gamma(x,y)\\ &=\underset{\phi\in\mathbb{R}^{N}}{\max}\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,N}\{\|x-y_{i}\|^{p}-\phi_{i}\}d\mu(x)+\sum_{i=1}^{N}\pi_{i}\phi_{i}.\end{split}$ (A.2)
@@
$W_{\epsilon,p}^{p}(\mu,\nu)=\underset{u,v\in\mathcal{C}(\mathbb{R}^{d})\times\mathbb{R}^{N}}{\max}\int_{\mathbb{R}^{d}}u(x)d\mu(x)+\sum_{i=1}^{N}\pi_{i}v_{i}-\epsilon\sum_{i=1}^{N}\pi_{i}\int_{\mathbb{R}^{d}}\exp\left(\frac{u(x)+v_{i}-\|x-y_{i}\|^{p}}{\epsilon}\right)d\mu(x)$ (A.3)
@@ A.2 The case of support points on the generalized diagonal $D_{N}$ and/or with null weights
For a functional of the form $(Y,\pi)\mapsto\mathbb{D}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$, where $\mathbb{D}$ is an optimal transport divergence, it is necessary to assume that the point cloud $Y$ lies outside of the generalized diagonal $D_{N}$, and that the weights are strictly positive in order to apply the duality formulations described above. Indeed, suppose that $y_{i}=y_{j}$, for some $i\neq j$, then the dual formulation of the problem should include the constraint $\phi_{i}=\phi_{j}$, as both $\phi_{i}$ and $\phi_{j}$ represent the evaluation of a continuous potential in the same argument. In other words, we would have $\phi_{i}=\phi(y_{i})=\phi(y_{j})=\phi_{j}$. While this may seem like a minor technicality that can often be overlooked, it cannot be omitted in our analysis. This is particularly relevant in the case of the sliced Wasserstein distance, where two distinct points in $\mathbb{R}^{d}$ may project onto the same point on a line. Similarly, it is important to ensure that the weights of the discrete measure remain in the interior of the probability simplex $\mathrm{int}(\Delta_{N})$.
@@
$W_{p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=W_{p}\left(\mu,\sum_{i=1}^{M}\tilde{\pi}_{i}\delta_{\tilde{y}_{i}}\right)$
and
$W_{\epsilon,p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=W_{\epsilon,p}\left(\mu,\sum_{i=1}^{M}\tilde{\pi}_{i}\delta_{\tilde{y}_{i}}\right).$
@@ A.3 Semi-duality formulation for semi-discrete sliced divergences
Starting from the definition of the sliced distances in (2.5), the dual formulation (A.2) can be applied to each one-dimensional distance $W_{p}(P_{\theta\sharp}\mu,P_{\theta\sharp}\nu),\>\theta\in\mathbb{S}^{d-1}$, addressing the ties in the support point of the measures $P_{\theta\sharp}\nu$. Let $(Y,\pi)\in(\mathbb{R}^{d})^{N}\setminus D_{N}\times\mathrm{int}(\Delta_{N})$, then, for each $\theta\in\mathbb{S}^{d-1}$ there exists an integer $M:=M(\theta)\in[\![1,N]\!]$ such that $\langle y_{i}-y_{j},\theta\rangle\neq 0$ and $\pi_{i}>0$ for all $i,j\in[\![1,M(\theta)]\!],\>i\neq j$. Up to a reordering of components, we can write $\pi_{i}$ as $\pi_{i}^{\theta}:=\pi_{i}+\pi_{j}$ whenever $\langle y_{i}-y_{j},\theta\rangle=0$ for all $i,j\in[\![1,N]\!],\>i\neq j$. Following this construction, we obtain a probability measure $\nu^{\theta}:=\sum_{i=1}^{M(\theta)}\pi_{i}^{\theta}\delta_{\langle y_{i},\theta\rangle}$ with no ties in its support points. Then, the Kantorovich duality formulation (A.2) leads to the following expressions of the sliced distances.
$SW_{p}^{p}(\mu,\nu)=\int_{\mathbb{S}^{d-1}}\left\{\underset{w\in\mathbb{R}^{M(\theta)}}{\max}\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,M(\theta)}{\min}\left\{|\langle x-y_{i},\theta\rangle|^{p}-w_{i}\right\}d\mu(x)+\sum_{i=1}^{M(\theta)}\pi_{i}w_{i}\right\}d\sigma(\theta)$ (A.5)
@@
where $\sigma$ denotes the uniform measure over the $d$-sphere $\mathbb{S}^{d-1}$. Note that actually, one can replace $M(\theta)$ by $N$ in the maximization over the variable $w$ in (A.5), without taking into account any dependency on $\theta$. Indeed, we can write
$M(\theta):=\mathrm{card}\left(\left\{i\in[\![1,N]\!]\,:\,\langle y_{i}-y_{j},\theta\rangle\neq 0,\>\forall j\neq i\right\}\right).$
We thus have $M(\theta)<N$ if and only if there exist indices $i\neq j$ such that $\langle y_{i}-y_{j},\theta\rangle=0$ , meaning that $\theta$ lies in
$\underset{i\neq j}{\bigcup_{i,j\in[\![1,N]\!]}}\left\{\theta\in\mathbb{S}^{d-1}\,:\,\langle y_{i}-y_{j},\theta\rangle=0\right\}.$
Since $Y\in(\mathbb{R}^{d})^{N}\setminus D_{N}$ this is a finite union of hyperplanes intersected with the sphere, which is $\sigma$-negligible. It follows that $M(\theta)=N$ for almost every $\theta\in\mathbb{S}^{d-1}$.

### v3

@@ 4.3.2 Dependency in $N$ in the estimation error
There is an $O\left(N^{-2/d}\right)$ factor between the lower bound in (4.1) and the upper bound in (3.1). This raises the question of the optimal exponent in $N$. We exhibit a situation where the lower bound is tight.
Setting $N=n$, $L=1$ and $\pi=(1/n,\ldots,1/n)$, the upper bound (3.1) of Corollary 3.4 remains constant as $n$ increases. However, we will argue that the error actually decreases like $O(n^{-2/d})$. Indeed, in this setting, the minimizer of $\mu\mapsto W_{2}^{2}(\mu_{n},\mu)$ over $\mathcal{M}_{1}^{N}(\mathbb{R}^{d})$, is clearly $\mu_{n}$ and thus the estimation error writes
$\mathbb{E}\left[W_{2}^{2}(\mu,\mu_{n})-\min_{\nu\in\mathcal{M}_{1}^{n}(\mathbb{B}_{R})}W_{2}^{2}(\mu,\nu)\right]\leq\mathbb{E}\left[W_{2}^{2}(\mu,\mu_{n})\right]=O(n^{-2/d}),$
see fournier2015rate. Hence Corollary 3.4 is not optimal in this particular case. The above argument suggests that the optimal exponent in $N$ is related to estimates of the quantity $W_{2}(\mu,\mu_{n})$, a notoriously difficult problem in statistical optimal transport (chewi2024statistical, Section 2). The curse of dimensionality is in fact hidden in the sparsity level $N$, as remarked in (del2024central, Theorem 2.10).
@@ 4.3.3 Beyond the $W_{2}^{2}$ case
We expect the same conclusions for $\mathbb{D}=SW_{p}^{p}$ and $\text{max-}SW_{p}^{p}$ since the $1/\sqrt{n}$ rate is also known to be optimal for these divergences xu2022central. The bias of the Sinkhorn divergence $\mathbb{D}=W_{\epsilon,p}^{p}$ benefits from a more favorable $1/n$ convergence rate del2023improved suggesting that a sharper rate is attainable for barycenters as well.
@@ 4.4 K-means and constrained K-means
Choosing $\mathbb{D}=W_{2}^{2}$ and $L=1$, Corollary 3.4 provides error bounds for the celebrated K-means problems. Let $X_{1},\ldots,X_{n}$ be $n$ i.i.d samples with distribution $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R})$. It is well known minimizing the K-means functional is actually an optimal quantization problem for the empirical measure $\mu_{n}$ (see Appendix D.4):
@@
Corollary 3.4 then yields a convergence rate of order $O(\sqrt{N/n})$ for the K-means problem, which yields an improvement by a factor $\sqrt{\log(N)}$ compared to (gyorfi2002principles, Section 4.4). The Wasserstein functional formulation is closer to the measure approximation view point. It also allows to consider the constrained K-means problem ng2000note, fixing the mass associated to each centroid with a weight vector $\bar{\pi}\in\Delta_{N}$ and minimizing the functional $Y\mapsto W_{2}^{2}\left(\mu_{n},\sum_{i=1}^{N}\bar{\pi}_{i}\delta_{y_{i}}\right)$. For uniform weights $\bar{\pi}_{i}=1/N$ for all $i\in[\![1,N]\!]$, this is the optimal uniform quantization problem to which Corollary 3.4 apply as well.
@@ 5 Numerical experiments
We illustrate [^1] The code to reproduce the experiments is available in Python at https://github.com/LeoPtls. Corollary 3.4 for $\mathbb{D}=W_{2}^{2}$ and $W_{\epsilon,2}^{2}$.
Dependency in $n$
We consider $L=3$ target Gaussian distributions in dimension $d=2$ with random means and covariances. We compute the $W_{2}^{2}$ and $W_{\epsilon,2}^{2}$ ($\epsilon=1$) empirical barycenter for $N=10$, $N=50$ and for $n$ ranging between $1000$ and $10000$ using the POT library flamary2021pot. Figure 1 represents the estimation error in (3.1) as $n$ increases, matching the rate in Corollary 3.4. The presence of hollows in the curve for $\mathbb{D}=W_{\epsilon,p}^{p}$ is due to numerical instabilities.
Figure 1: The graphs show the evolution of the estimation error (in log-scale) as the sample size $n$ increases. Each experiments are performed $40$ times. The mean value of the estimation errors is plotted along with the $95\%$ confidence interval. The theoretical upper bound $n\mapsto 1/\sqrt{n}$ is plotted in red.
Dependency in NN
We consider the same setting as above in dimension $d=6$ for which the theoretical rates predict larger errors as $N$ increases. We fix the sample size $n=500$ and consider $N$ ranging between $10$ and $500$. Figure 2 shows the estimation error as a function of $N$ together with the upper bound of Corollary 3.4 and lower bound of Lemma 4.1 for $\mathbb{D}=W_{2}^{2}$. This experiment suggests that the dependency in $N$ given by the worst case scenario in Lemma 4.1 for $W_{2}^{2}$ matches practical observations on the mild gaussian distributions setting.
Figure 2: The graphs show the estimation error (in log-scale) as the support cardinality NN increases, for $n=500$. Each experiments are performed $10$ times. The mean value of the estimation errors is plotted along with the $95\%$ confidence interval. The theoretical bounds (Corollary 3.4 and Lemma 4.1) are plotted in red.
@@ 6 Conclusion
The expected convergence rate of the empirical sparse optimal transport barycenters is shown to be of order $O\left(\sqrt{N/n}\right)$ for several types of optimal transport divergences. The rate is min-max optimal for the Wasserstein distance $\mathbb{D}=W_{2}^{2}$. Investigating optimality for sliced distances $\mathbb{D}=SW_{p}^{p},\>\text{max-}SW_{p}^{p}$ and Sinkhorn divergences $\mathbb{D}=W_{\epsilon,p}^{p}$ is left for future work.
@@ Appendix A Semi-dual formulation of optimal transport
The Kantorovich dual formulation of the Wasserstein distances $W_{p}$ can be recast as a semi-dual problem using the $c$-transform. Similarly, the Sinkhorn divergences $W_{\epsilon,p}^{p}$ admit a semi-dual formulation based on the smoothed $c$-transform. In the semi-discrete setting (i.e. between a continuous and a discrete measure), these formulations turn an infinite-dimensional program (corresponding to the primal formulations (2.1) and (2.3)) into a finite-dimensional one, which is a key point in our proofs. In this section, we present these semi-dual formulations in the semi-discrete setting, with a particular focus on the case where the discrete measure has zero weights or repeated atoms. Throughout this section, we consider a measure $\mu\in\mathcal{M}_{1}(\mathbb{R}^{d})$.
@@ A.1 Semi-dual formulation outside the generalized diagonal
Let us then define $\nu:=\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}$ a discrete measure supported on $Y:=(y_{1},\ldots,y_{N})\in(\mathbb{R}^{d})^{N}\setminus D_{N}$ and strictly positive weights $\pi\in\mathrm{int}(\Delta_{N})$. For a bounded continuous function $\phi\in\mathcal{C}_{b}(\mathbb{R}^{d})$ , let us also denote $\phi^{c}\in\mathcal{C}_{b}(\mathbb{R}^{d})$ its c-transform, defined in our case for all $x\in\mathbb{R}^{d}$ as $\phi^{c}(x)=\min_{i=1,...,N}(\|x-y_{i}\|^{p}-\phi(y_{i}))$ (book_santambrogio, Section 1.3). Then by Kantorovich’s duality ((villani_opt_old_new, Theorem 5.10)) and $c$-convexity of the optimal potentials we have the following formulation for the $p$-Wasserstein distance between $\mu$ and $\nu$:
$W_{p}^{p}(\mu,\nu)=\underset{\phi\in\mathbb{R}^{N}}{\max}\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,N}\{\|x-y_{i}\|^{p}-\phi_{i}\}d\mu(x)+\sum_{i=1}^{N}\pi_{i}\phi_{i}.$ (A.2)
@@
$\begin{split}W_{\epsilon,p}^{p}(\mu,\nu)&=\underset{u,v\in\mathcal{C}(\mathbb{R}^{d})\times\mathbb{R}^{N}}{\max}\int_{\mathbb{R}^{d}}u(x)d\mu(x)+\sum_{i=1}^{N}\pi_{i}v_{i}\\ &\qquad-\epsilon\sum_{i=1}^{N}\pi_{i}\int_{\mathbb{R}^{d}}\exp\left(\frac{u(x)+v_{i}-\|x-y_{i}\|^{p}}{\epsilon}\right)d\mu(x)\end{split}$ (A.3)
@@ A.2 The case of support points on the generalized diagonal $D_{N}$ and/or with null weights
As previously mentioned, it is mandatory that $(Y,\pi)\in\big(\mathbb{R}^{d}\big)^{N}\setminus D_{N}\times\mathrm{int}(\Delta_{N})$ for a functional of the form $(Y,\pi)\mapsto\mathbb{D}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$ to admit the duality formulations described above. Indeed, suppose that $y_{i}=y_{j}$, for some $i\neq j$, then the dual formulation of the problem should include the constraint $\phi_{i}=\phi_{j}$, as both $\phi_{i}$ and $\phi_{j}$ represent the evaluation of a continuous potential in the same argument. In other words, we would have $\phi_{i}=\phi(y_{i})=\phi(y_{j})=\phi_{j}$. While this may seem like a minor technicality that can often be overlooked, it cannot be omitted in our analysis. This is particularly relevant in the case of the sliced Wasserstein distance, where two distinct points in $\mathbb{R}^{d}$ may project onto the same point on a line. Similarly, it is important to ensure that the weights of the discrete measure remain in the interior of the probability simplex $\mathrm{int}(\Delta_{N})$.
@@
$W_{p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=W_{p}\left(\mu,\sum_{i=1}^{M}\tilde{\pi}_{i}\delta_{\tilde{y}_{i}}\right)\quad\text{and}\quad W_{\epsilon,p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=W_{\epsilon,p}\left(\mu,\sum_{i=1}^{M}\tilde{\pi}_{i}\delta_{\tilde{y}_{i}}\right).$
@@ A.3 Semi-duality formulation for semi-discrete sliced divergences
Starting from the definition of the sliced distances in (2.4), the dual formulation (A.2) can be applied to each one-dimensional distance $W_{p}(P_{\theta\sharp}\mu,P_{\theta\sharp}\nu),\>\theta\in\mathbb{S}^{d-1}$, addressing the ties in the support point of the measures $P_{\theta\sharp}\nu$. Let $(Y,\pi)\in(\mathbb{R}^{d})^{N}\setminus D_{N}\times\mathrm{int}(\Delta_{N})$, then, for each $\theta\in\mathbb{S}^{d-1}$ there exists an integer $M:=M(\theta)\in[\![1,N]\!]$ such that $\langle y_{i}-y_{j},\theta\rangle\neq 0$ and $\pi_{i}>0$ for all $i,j\in[\![1,M(\theta)]\!],\>i\neq j$. Up to a reordering of components, we can write $\pi_{i}$ as $\pi_{i}^{\theta}:=\pi_{i}+\pi_{j}$ whenever $\langle y_{i}-y_{j},\theta\rangle=0$ for all $i,j\in[\![1,N]\!],\>i\neq j$. Following this construction, we obtain a probability measure $\nu^{\theta}:=\sum_{i=1}^{M(\theta)}\pi_{i}^{\theta}\delta_{\langle y_{i},\theta\rangle}$ with no ties in its support points. Then, the Kantorovich duality formulation (A.2) leads to the following expressions of the sliced distances :
$SW_{p}^{p}(\mu,\nu)=\int_{\mathbb{S}^{d-1}}\left[\underset{w\in\mathbb{R}^{M(\theta)}}{\max}\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,M(\theta)}{\min}\left\{|\langle x-y_{i},\theta\rangle|^{p}-w_{i}\right\}d\mu(x)+\sum_{i=1}^{M(\theta)}\pi_{i}w_{i}\right]d\sigma(\theta)$ (A.5)
@@
where $\sigma$ denotes the uniform measure over the $d$-sphere $\mathbb{S}^{d-1}$. Note that actually, one can replace $M(\theta)$ by $N$ in the maximization over the variable $w$ in (A.5), without taking into account any dependency on $\theta$. Indeed, let $M(\theta):=\mathrm{card}\left(\left\{i\in[\![1,N]\!]\,:\,\langle y_{i}-y_{j},\theta\rangle\neq 0,\>\forall j\neq i\right\}\right).$ We thus have $M(\theta)<N$ if and only if there exist indices $i\neq j$ such that $\langle y_{i}-y_{j},\theta\rangle=0$, meaning that $\theta$ lies in a union of sets of the form $\left\{\theta\in\mathbb{S}^{d-1}\,:\,\langle y_{i}-y_{j},\theta\rangle=0\right\},$ for $i,j\in[\![1,N]\!]$ and $i\neq j$. Since $Y\in(\mathbb{R}^{d})^{N}\setminus D_{N}$ this is a finite union of hyperplanes intersected with the sphere, which is $\sigma$-negligible. It follows that $M(\theta)=N$ for almost every $\theta\in\mathbb{S}^{d-1}$.